OpenStudy (richgall____):

The hypotenuse of a right triangle is 12, and one of the legs is 4. Find the length of the missing leg to the nearest hundredth.

pooja195 (pooja195):

We know that the hypotenuse is the C value and 4 can be either A or B We need to use the Pythagorean theorem to figure out the missing length \[\huge~\rm\bf~a^2+b^2=c^2 \] When we plug in the numbers we get this: \[\huge~\rm~4^2+b^2=12^2\] Try solving for b :-)

OpenStudy (richgall____):

4^2=16 12^2=144 so know i just have to find the amount that goes between those 2?

pooja195 (pooja195):

Not quite ok now that you found those numbers \[\huge~\rm~\bf~16+b^2=144\] We want to isolate b like get it by itself Any ideas on what we should do?

OpenStudy (richgall____):

im not quite sure im just now really getting to know this topic

pooja195 (pooja195):

I see since we want to get b by it self we would subtract 16 from both sides \[\huge~\rm~\bf~b^2+16−16=144−16\] what are we left with?

OpenStudy (richgall____):

doesnt the first part cancel out since you're subtracting 16

pooja195 (pooja195):

Yes :) now you have 144=16=?

OpenStudy (richgall____):

128

pooja195 (pooja195):

Good \[\huge~\rm~b^2=128\] To get b by itself we need to square root both sides \[\huge~\rm~\bf~\sqrt{b^2}=\sqrt{128}\] Now we have \[\huge~\rm~\bf~b=\sqrt{128}\ since your question asks for a decimal form what is the square root of 128?

OpenStudy (richgall____):

11.31 thats if its rounded to the nearest hundredth

pooja195 (pooja195):

That is correct :-)

OpenStudy (richgall____):

thanks

pooja195 (pooja195):

yw ^_^